OpenStudy (anonymous):

one eighth of a pizza sold for $.95. (Cost of the whole pizza)

OpenStudy (seaturtle113):

multiply .95 by 8.

OpenStudy (anonymous):

i have to write an equation that represents the given information

OpenStudy (seaturtle113):

Which is $7.60

OpenStudy (seaturtle113):

Okay.. 1/8 = .95 therefore .95 * 8 =8/8 I think that's correct, not 100%, though

OpenStudy (anonymous):

no, like for example: A book store has sold all but 12 of the dictionaries in a shipment of 120 dictionaries. (Number of dictionaries sold) so that would be x+4=30

OpenStudy (anonymous):

i mean x+12=120

OpenStudy (seaturtle113):

Ahh... I don't know then, sorry. :/

OpenStudy (anonymous):

ok then thx

OpenStudy (anonymous):

If you need an equation, it has to start with a variable, right? So what's the variable that we are interested in?

OpenStudy (anonymous):

is it 1/8(.95)=x

OpenStudy (seaturtle113):

I think so!!

OpenStudy (anonymous):

ok thnx